The stockholders’ equity section of the balance sheet for Mann Equipment Co. at December 31, Year 1, is as follows:

Stockholders’ Equity
Paid-in capital
Preferred stock, ? par value, 4% cumulative,
120,000 shares authorized, 42,000 shares issued and outstanding
$ 420,000
Common stock, $20 stated value, 170,000 shares
authorized, 42,000 shares issued and outstanding
840,000
Paid-in capital in excess of par—Preferred 32,000
Paid-in capital in excess of stated value—Common 105,000
Total paid-in capital 1,397,000
Retained earnings 270,000
Total stockholders’ equity $ 1,667,000


Note: The market value per share of the common stock is $44, and the market value per share of the preferred stock is $14.

Required
a. What is the par value per share of the preferred stock?
b. What is the dividend per share on the preferred stock? (Round your answer to 2 decimal places.)
c. What was the average issue price per share (price for which the stock was issued) of the common stock? (Round your answer to 2 decimal places.)
e. If Mann declares a 2-for-1 stock split on the common stock, how many shares will be outstanding after the split? What amount will be transferred from the retained earnings account because of the stock split? Theoretically, what will be the market price of the common stock immediately after the stock split?
Answer #1

a.

Par value per share of preferred stock = Preferred stock/Number of preferred shares

= 420,000/42,000

= $10

b.

Dividend per share on the Preferred stock = Par value per preferred share x Dividend rate

= 10 x 4%

= $0.40

c.

Average issue price per share of common stock = (Common stock + Paid in capital in excess of par common)/Number of common shares issued

= (840,000+105,000)/42,000

= 945,000/42,000

= $22.50

e.

After 2- for - 1 stock split, the number of outstanding shares of common stock will be = 42,000 x 2

= 84,000

Retained earnings will not be effected dud to stock split.

Market price per share after stock split = 44/2

= $22

Kindly comment if you need further assistance.

Thanks‼!